Silver Company makes a product that is very popular as a Mother’s Day gift. Thus, peak sales occur in May of each year, as shown in the company’s sales budget for the second quarter given below:

April May June Total
Budgeted sales (all on account) $440,000 $640,000 $220,000 $1,300,000

From past experience, the company has learned that 30% of a month’s sales are collected in the month of sale, another 60% are collected in the month following sale, and the remaining 10% are collected in the second month following sale. Bad debts are negligible and can be ignored. February sales totaled $370,000, and March sales totaled $400,000.

Required:

1. Prepare a schedule of expected cash collections from sales, by month and in total, for the second quarter.

2. What is the accounts receivable balance on June 30th?
Answer #1

Answer:

Silver Company
Schedule of expected cash collections from sale
Particulars April May June Total
February sales $370000*10%=$37000 0 0                 37,000
March sales $400000*60%=$240000 $400000*10%=$40000 0              2,80,000
April sales $440000*30%=$132000 $440000*60%=$264000 $440000*10%=$44000              4,40,000
May sales 0 $640000*30%=$192000 $640000*60%=$384000              5,76,000
June sales 0 0 $220000*30%=$66000                 66,000
Total cash collections                            4,09,000                            4,96,000                               4,94,000            13,99,000
2)The accounts receivable balance on june 30th = ($640000*10%)+($220000*70%)
=$64000+$154000=$218000